# On the decoherence of Majorana zero modes mediated by gapless fermions

### v1

@@ II.2 Lindblad equation for the Majorana zero modes
The derivation of the Lindblad equation from the microscopic Hamiltonian (2) is carried out in App. A. It adheres quite closely to textbook derivations of master equations from microscopic models [29], with the specificity that it is carried out entirely in terms of Majorana operators. A similar weak-coupling derivation for a free fermionic model was also provided in Ref. [28]. The properties of $\Lambda$ and $J$ are also derived by construction in App. A.
@@ II.3.3 Non-locality
The long-range couplings between Majorana zero modes cannot be established instantaneously. Indeed, the validity of the Lindblad equation (3) kicks in only after a time long enough such that correlations between distant Majorana zero modes can be established via the reservoir. For the Lindblad description to be valid, this time has to be smaller than the inverse tunneling rate between a Majorana zero modes and the reservoir. Therefore, the larger the distance between Majorana zero modes, the weaker the tunnel couplings need to be in order for the Lindblad equation to be valid. This aspect is further addressed in the exact numerical calculations of Sec. V, and in the derivation of App. A.
II.3.4 Steady states and spectral gap
@@ II.4 Time evolution of Gaussian states
While the Lindblad equation is valid for an arbitrary initial state of the Majorana zero modes, we will from now on assume that the initial state is Gaussian. Such a state $\rho$ is completely specified by the covariance matrix [37, 38, 39]:
@@ III Occurrence of slow parity decay
with $r^{2}=(\pi J_{12}-\Lambda_{12})/(\pi J_{12}+\Lambda_{12})$.
@@ V Comparison to numerics
The critical Kitaev chain has a gapless spectrum, with excitation energies $\zeta(k)=2h\left|\sin(k/2)\right|$. The couplings $\Lambda_{ij}$ and $J_{ij}$ can be computed analytically from a diagonalization of the critical Kitaev chain, as done in App. C. Letting $d_{ij}=\left|n_{i}-n_{j}\right|$ be the distance between two Majorana zero modes along the chain (in units of the lattice spacing), the results are
@@ V.1 Setup and methods
Figure 6: Time evolution of various parity expectation values for the four-Majorana, two-reservoir system of Fig. 5, for even distances $d$ between Majorana zero modes. The blue and red curves correspond to X-type and Z-type initial conditions respectively. The rightmost panel shows the time evolution of the total parity, which is the same for both initial conditions. The dashed curves are the Lindblad predictions, which for this model are independent of $d$, while the solid lines represent numerical results for different values of $d$, increasing in steps of 10 from $d=0$ to $d=100$ from the darker to the lighter lines.
@@ V.2 Numerical results
We interpret these results as follows. When the temperature of the reservoir becomes larger than the Lamb shifts, we recover the fully mixed steady state corresponding to an exact degeneracy of Majorana zero modes, as predicted by the Lindblad equation. On the other hand, if the temperature of the reservoir is smaller than the Lamb shift, the steady state (which is not by fully accessible in our simulation due to finite size-effects) deviates from the Lindblad prediction.
@@ VI Conclusions
Our findings leave some immediate open questions. Can the Lindblad equation be refined, perhaps with a better time coarse-graining, to capture the time delay occurring before non-local couplings are established? Can the spectral properties of the Lindblad equation for the free-electron model be described with random-matrix theory? Can the open-system dynamics reveal a difference between unpaired Majorana zero modes and local zero-energy Andreev bound state [42]? And, perhaps most importantly, can the non-local decoherence effect be observed in near-term experiments with Majorana devices? We leave the exploration of these questions for further work.
@@ Appendix D Comments on the reservoir correlation function
Here $x_{i}$ and $x_{j}$ denote the positions of the $i$-th and $j$-th Majorana modes respectively. After putting in the above expressions in Eq. 70 and simplifying for zero temperature ($\beta\rightarrow\infty$), we obtain,
@@
This is an oscillatory function with a power law ($\propto 1/\sqrt{t}$) decay. However, the equation of motion for $\rho$ contains only integrals of this correlation function over time. These indeed have a dominant contribution for small times, after which they merely oscillate about a mean value. This lends validity to the Born-Markov approximation we invoked to derive the effective Lindblad equation 3. We also comment that in presence of potential disorder, the correlation function shall also contain an exponential decay, with a time scale proportional to $1/W$, where $W$ is the disorder strength. Provided this scale is much shorter than the timescales corresponding to the system-reservoir coupling, again Markov approximation may be justified.

### v2

# Decoherence of Majorana zero modes mediated by gapless fermions

@@ II.2 Lindblad equation for the Majorana zero modes
The derivation of the Lindblad equation from the microscopic Hamiltonian (2) is carried out in App. A. It adheres quite closely to textbook derivations of master equations from microscopic models [11], with the specificity that it is carried out entirely in terms of Majorana operators. A similar weak-coupling derivation for a free fermionic model was also provided in Ref. [13]. The properties of $\Lambda$ and $J$ are derived by construction in App. A.
@@ II.3.3 Non-locality
II.3.4 Validity of the weak coupling approximation
The Lindblad master equation is based on the Born-Markov approximation, thus on a perturbative expansion valid for weak couplings between the Majorana zero modes and the bath. For the approximation to be valid, to begin with the couplings $W_{in}$ in Eq. (2) must be much smaller than the topological superconducting gap in the system hosting the Majorana zero modes. Otherwise, the Hamiltonian (2) would not be a good starting point to model the system-reservoir interaction. Furthermore, as emphasized in recent literature on open quantum systems [37, 35], the validity of the Lindblad equation rests on the comparison between the bath correlation time on one hand and the coupling strength between the system and the reservoir on the other hand.
In particular, it is clear that the long-range couplings between Majorana zero modes appearing in the Lindblad equation cannot be established instantaneously. Thus, we must require that the inverse tunneling rate be much longer than the time it takes to establish a correlation between distant Majorana zero modes via the reservoir. There are, indeed, multiple two-point bath correlation functions that must be evaluated in order to derive the Lindblad equation - one for each non-local term in the dissipator, as detailed in App. A.
Because the correlation time will increase with the distance between Majorana zero modes, broadly speaking the larger the distance, the weaker the tunnel couplings need to be in order for the Lindblad equation to be valid. Increasing the coupling strengths between the Majorana zero modes and the reservoir or the distance between Majorana zero modes, non-Markovian effects become more relevant, causing larger deviations from the predictions of the Lindblad equation. These aspects are further addressed in the exact numerical calculations of Sec. V, see in particular Fig. 7, and in the derivation of App. A.
II.3.5 Steady states and spectral gap
@@ II.4 Time evolution of Gaussian states
While the Lindblad equation is valid for an arbitrary initial state of the Majorana zero modes, we will from now on assume that the initial state is Gaussian. Such a state $\rho$ is completely specified by the covariance matrix [46, 10, 9, 8]:
@@ III Occurrence of slow parity decay
with $r^{2}=(\pi J_{12}+\Lambda_{12})/(\pi J_{12}-\Lambda_{12})$.
@@ V Comparison to numerics
The critical Kitaev chain has a gapless spectrum, with excitation energies $\zeta(k)=2h\left|\sin(k/2)\right|$. The couplings $\Lambda_{ij}$ and $J_{ij}$ can be computed analytically from a diagonalization of the critical Kitaev chain, as done in App. C. Letting $d_{ij}$ be the distance (in units of the lattice spacing) between the two sites of the chain to which $a_{i}$ and $a_{j}$ are coupled to, the results are
@@ V.1 Setup and methods
Figure 6: Time evolution of various parity expectation values for the four-Majorana, two-reservoir system of Fig. 5, for even distances $d$ between Majorana zero modes. The blue and red curves correspond to X-type and Z-type initial conditions respectively. The rightmost panel shows the time evolution of the total parity, which is the same for both initial conditions. The dashed curves are the Lindblad predictions, which for this model are independent of $d$, while the solid lines represent numerical results for different values of $d$, increasing in steps of 10 from $d=0$ to $d=100$ from the darker to the lighter lines. The dashed vertical line in the leftmost panel indicates the time slice at which the further results presented in Fig. 7 are computed.
Figure 7: Evolution of the parity expectation value $M_{14}^{X}$ computed at $t=500$ as a function of the distance $d$ between the Majorana zero modes. Different curves correspond to increasing values of the coupling strength $w$, from $0.1$ to $0.8$ in steps of $0.1$. The plot shows more systematically the decrease in the expectation value, computed at the plateau reached in Fig. 6 before the onset of the finite-size revival, when increasing distance or coupling strength.
@@ V.2 Numerical results
This effect is illustrated in Fig. 7, where we show the evolution of the plateau reached in Fig. 6 before the onset of finite-size effects. When increasing the distance between Majorana zero modes or the coupling strength between the Majorana zero modes and the reservoir, the numerical calculations deviate ever more from the steady-state value predicted by the Lindblad with non-local terms. Indeed, when correlations between distant zero modes cannot be established quickly enough, the non-local effects vanish and the predictions for a disconnected reservoir are re-established.
@@
We interpret these results as follows. When the temperature of the reservoir becomes larger than the Lamb shifts, we recover the fully mixed steady state corresponding to an exact degeneracy of Majorana zero modes, as predicted by the Lindblad equation. On the other hand, if the temperature of the reservoir is smaller than the Lamb shift, the steady state (which is not fully accessible in our simulation due to finite size-effects) deviates from the Lindblad prediction.
@@ VI Conclusions
Although our modeling did not include the fermion parity measurement apparatus explicitly, we emphasize that the protocol described in Sec. III to isolate the non-local decoherence effect could be experimentally implemented e.g. using tetron devices [25, 32]. In this setting, a two-dimensional electron gas could be used as an intentional fermionic reservoir with tunable couplings, in order to test the dissipative dynamics of Majorana zero modes. Conceptually, the presence of two separate baths is not strictly necessary: it would be sufficient to couple two out of four Majorana zero modes to a common reservoir.
Our findings leave some immediate open questions. Can the Lindblad equation be refined, perhaps with a better time coarse-graining, to capture the time delay occurring before non-local couplings are established? Can the spectral properties of the Lindblad equation for the free-electron model be described with random-matrix theory? Can the open-system dynamics reveal a difference between unpaired Majorana zero modes and local zero-energy Andreev bound state [33]? We leave the exploration of these questions for further work.
@@
Data availability
All data generated and code used in this work are available on Zenodo [47].
@@ Appendix D Comments on the reservoir correlation function
Here $x_{i}$ and $x_{j}$ denote the positions of the $i$-th and $j$-th Majorana modes respectively. After putting in the above expressions in Eq. 70 and simplifying for zero temperature ($\beta\rightarrow\infty$), we obtain
@@
This is an oscillatory function with a power law ($\propto 1/\sqrt{t}$) decay. However, the equation of motion for $\rho$ contains only integrals of this correlation function over time. These indeed have a dominant contribution for small times, after which they merely oscillate about a mean value.
This lends validity to the Born-Markov approximation we invoked to derive the effective Lindblad equation (3). We also comment that in presence of potential disorder, the correlation function shall also contain an exponential decay, with a time scale proportional to $1/W$, where $W$ is the disorder strength. Provided this scale is much shorter than the timescales corresponding to the system-reservoir coupling, again Markov approximation may be justified.